This is a very big test for me and I really need help here Please

in 1990 the life expectency of males in a certain country was 64.2 yrs. In 1995 it was 66.9yrs. Let E represent the life expectency in the yr t and let t represent the number of yrs since 1990.
the linear functionE(t) that fits the data is
use the function to predict the life expectency of males in 2004

Answer by scott8148(6628)

in 5 years, the life expectancy increased by 2.7 years ___ it increases .54 yr per yr ___ this is the slope (m) of the linear function

E(t) = .54t + b ___ to find the constant term (intercept), plug in a data point

E(5) = .54(5) + b ___ 66.9 = 2.7 + b ___ 64.2 = b ___ this makes sense because 1990 was the base year

E(t) = .54(t) + 64.2

2004 is 14 years since 1990 ___ substitute 14 for t to predict the life expectancy